# Differentially Private Relational Learning with Entity-level Privacy Guarantees

### v1

@@ 4.2 Coupled Sampling and Its Privacy Amplification
Here coupling refers to the dependency of $\mathcal{B}^{(2)}$ on $\mathcal{B}^{(1)}$. Coupled sampling is a general framework that covers most relational learning sampling settings. For instance, a common practice is to first sample some positive edges and “perturb” one end of these positive edges to form negative edges (bordes2013translating, ). In this case, $D^{(1)}=\mathcal{E},D^{(2)}=\mathcal{E}^{c}$ (non-edges), and: (1) $p(\mathcal{B}^{(1)}|D^{(1)})=p(e^{+}_{1},e^{+}_{2},...|\mathcal{E})$ randomly sample a few positive edges from $\mathcal{E}$ (e.g., Poisson sampling); (2) to sample negative edges $\mathcal{B}^{(2)}_{i}=(e^{-}_{i,1},...,e^{-}_{i,{k_{\mathrm{neg}}}})$, first randomly sample one end of $e_{i}^{+}$, denoted by $\tilde{u}_{i}$, and then sample $k_{\mathrm{neg}}$ negative edges $\mathcal{B}^{(2)}_{i}$ from $\{(\tilde{u}_{i},v):(\tilde{u}_{i},v)\in\mathcal{E}^{c}\}$, i.e., non-edges localized at $\tilde{u}_{i}$; (3) the resulting mini-batch is =$\mathcal{B}=\{(e_{i}^{+},e^{-}_{i,1},...,e^{-}_{i,{k_{\mathrm{neg}}}}):i=1,...,|\mathcal{B}^{(1)}|\}$.
Coupled sampling extends the existing privacy amplification framework by generalizing the sampling process to a multi-step, dependent procedure, where the flexible and potentially intricate dependencies between steps pose new analytical challenges. In the absence of such dependencies, that is, $\mathcal{B}^{(2)}\sim q(\cdot\mid D^{(2)})$, the contributions of each sampling step can be decomposed, allowing existing privacy amplification results to be adapted: in this decoupled case, the mini-batch =$\mathcal{B}=\{(\mathcal{B}^{(1)}_{i},\mathcal{B}^{(2)}_{i})\}_{i=1,2,\ldots}$ can be viewed as a single-step sampling from the Cartesian product dataset $D^{(1)}\times D^{(2)}$. However, when dependencies are introduced, this separation no longer holds.
@@
    $w\leftarrow\text{a random end in }e_{i}^{+}$, and $e_{i,j}^{-}\leftarrow(w,v_{i,j})$triangleright pair $w$ with the sampled node $v_{i,j}$
Return =$\mathcal{B}=\{T_{1},T_{2},...,T_{b}\}=\{(e_{i}^{+},e_{i,1}^{-},...,e_{i,k_{\mathrm{neg}}}^{-})\}_{i=1,...,b}$
@@
Initialize Randomly drop edges from $\mathcal{E}$, so that maximal node degree is capped by $K$ and obtain $\bar{\mathcal{E}}$ (daigavane2021node, , Algorithm 1). Let positive sampling rate $\gamma\leftarrow b/|\bar{\mathcal{E}}|$$\triangleright$ Node degree capping
@@
    positive edges $E^{+}$leftarrow independently choosing each positive relation from $\bar{\mathcal{E}}$ with probability $\gamma$.
@@
    $\bar{\mathbf{g}}_{t}\leftarrow\text{FREQ-CLIP}(\mathcal{B}_{t},\{\mathbf{g}_{t}(T_{1}),\mathbf{g}_{t}(T_{2}),...\},C)$triangleright Adaptive clipping Algorithm 1
@@
In the context of relational learning, cardinality-dependent sampling can be realized by choosing negative sampling methods that are independent of the contents of the sampled positive edges. For instance, we adopt Poisson subsampling as positive sampling, and adopt Algorithm 2 for negative sampling, which first randomly draws some nodes from the node set (independent of specific positive edges), and then constructs contrastive pairs by pairing one end of positive edges with the sampled nodes. Note that this pairing may occasionally produce negative edges that are actually true positives—a known issue in node pairing methods like in-batch sampling (you2020graph, ). However, this occurs with a negligible probability, especially given the sparsity of real-world graphs.
@@
where $\mathrm{Bin}(m,\gamma)$ stands for the Binomial distribution, effective sampling rate $\Gamma_{\ell}\coloneqq 1-(1-\gamma)^{K}(1-\frac{\ell\cdot k_{\mathrm{neg}}}{n})$ and $\Psi_{\alpha}(P\|Q)\coloneqq\mathbb{E}_{x\sim Q}(P(x)/Q(x))^{\alpha}$ for two distributions $P,Q$. The proof is deferred to Appendix B. $\Psi_{\alpha}$ here is a univariate integral and can be numerically calculated up to arbitrary precision (mironov2019r, ). Although Theorem 3 adopts $S^{(1)}$ as Poisson sampling and $S^{(2)}$ as sampling without replacement, the analysis and results can be extended to other combinations of sampling strategies. We leave these generalizations to the extended version of this work.
@@
Note that the bound Eq.(3) depends on the maximal node degree $K$. To mitigate the influence of high-degree nodes, in practice we cap each node’s degree at a desired constant $K$ by randomly sampling its neighbors, same as Algorithm 1 of (daigavane2021node, ). The effect of degree capping on utility is explored in Figure 3 in our later experiments.
@@ 5 Experimental Results
Per-iteration RDP Bounds. We evaluate the RDP bound $\varepsilon(\alpha)$ (Eq.(3)), shown in the left two subfigures of Figure 2, which characterizes the per-iteration privacy loss in DP-SGD. Our bound is compared against two baselines: (a) the naive Gaussian mechanism bound $\varepsilon_{\mathrm{Gaussian}}(\alpha)=\alpha/2\sigma^{2}$ (mironov2017renyi, ) without any amplification analyses; (b) the amplification bound with standard gradient clipping in place of the proposed adaptive clipping. The latter is also a coupled-sampling amplification bound we derived using an analysis similar to that of Theorem 3, and is provided in Appendix C.
DP Bounds over Composition. We also compare the accumulated privacy loss of applying multiple iterations of DP-SGD in terms of $(\varepsilon,\delta)$-DP, tracked by composition theorems of RDP (mironov2017renyi, ) and translating RDP to DP balle2020hypothesis . We choose privacy budget $\delta=1/m=0.2\times 10^{-6}$ and compare $\varepsilon_{\mathrm{DP}}(T)$ as a function of iteration $T$. The results of DP bound over composition are shown in the right two subfigures in Figure 2.
@@ 5.2 Private Relational Learning for Pretrained Text Encoders
Figure 3: Utility of different ##\##negative sample per positive $k_{\text{neg}}$, degree capping $K$, and noise multiplier $\sigma$ for zero-shot relation prediction. The legend AMAZ-sports means training on AMAZ-cloth and testing on AMAZ-sports, and similar for the legend AMAZ-cloth. In the rightmost figure, numbers along each line indicate the corresponding privacy parameter $\varepsilon_{\mathrm{DP}}$ at different noise levels.
Experimental Settings. We adopt two pre-trained language models BERT (devlin2018bert, ) and Llama2 (touvron2023llama, ) as the text encoders, and four text-attributed graphs from two subdomain pairs: two citation networks (MAG-CHN, MAG-USA) (sinha2015overview, ) and two co-purchase networks (AMZA-Sports, AMZA-Cloth) (mcauley2015image, ). The models are first privately fine-tuned on one network (e.g., MAG-CHN), and then its utility is evaluated on another same-domain network (e.g., MAG-USA). We report the overall privacy loss in terms of $(\varepsilon,\delta)$-DP, where $\delta$ is set to $1/|\mathcal{E}_{\text{train}}|$, the size of the relation set used for training after degree capping. We use the ranking metrics of top@1 precision (PREC@1) and mean reciprocal rank (MRR) to evaluate the relation prediction performance. The maximal node degree is capped by $K=5$ according to Algorithm 3. See Appendix E for more implementation details.
@@ Appendix B Proof of Theorem 3
Notation. We use the removal/insertion notation of DP: we have two neighboring graphs $G=(\mathcal{V},\mathcal{E}),G^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ where $\mathcal{V}=\{1,2,...,n\}$, $\mathcal{V}^{\prime}=\{1,2,...,n-1\}$ and =$\mathcal{E}=\mathcal{E}^{\prime}\cup\{(n,u_{i}):i=1,2,...,K\}$. Let $m=|\mathcal{E}|$. Let $\mathbb{X}=2^{\mathcal{E}},\mathbb{X}^{\prime}=2^{\mathcal{E}^{\prime}}$ be the all possible subset of edges that represents the space of positive samplings. Let $\mathbb{Y}=\{(y_{1},y_{2}...,y_{b\cdot k_{\mathrm{neg}}}):y_{i}\in V,b\in[m],y_{i}\neq y_{j}\}$ represent all possible ordered node set that represents the space of negative samplings. For simplicity, we will assume $k_{\mathrm{neg}}=1$, i.e., each positive sample is only paired with one negative sample. In the end we will show how to adapt to case $k_{\mathrm{neg}}>1$. An actual dataset we are computing on is $(x,y)$ for some $x\in\mathbb{X},y\in\mathbb{Y}$ or $(x^{\prime},y^{\prime})$ for some $x^{\prime}\in\mathbb{X}^{\prime},y\in\mathbb{Y}^{\prime}$ (again, $\mathbb{X}^{\prime},\mathbb{Y}^{\prime}$ are the sampling space of positive edges and ordered node set defined on neighboring graph $\mathcal{G}^{\prime}$). We denote $\mu_{x,y}$ as the output distribution of the mechanism acting on a fixed sampled dataset $(x,y)$, and denote $p,q$ as the mixture output distribution of the mechanism acting on a randomly sampled dataset $(x,y)$ and $(x^{\prime},y^{\prime})$ respectively. By definition, we have
@@
Jensen’s inequality by conditional coupling. The key idea is to apply Jensen’s inequality to $\Psi_{\alpha}(\cdot||\cdot)$ as $\Psi_{\alpha}$ is jointly convex [80]. To do so, we need to re-write $p,q$ in terms of balancing summation. we will generalize the conditional coupling technique [54] to our case. In the following we will show how we can achieve this in a high-level view. Consider the conditional distributions $\omega(x|A_{i})$, where $A_{i}$ represents the event “$x$ has $i$ removed edges (i.e., the edges involve the removed node $n$)”, and conditional distributions $\omega(y|B_{j},x)$, $j=0,1$, where $B_{0}$($B_{1}$) represents the event “$[y]_{S(x)}$ has (no) removed node $n$”, where $S(x)$ is a subset of indices defined by $S(x)\equiv\{\ell\in\{1,2,...,|x|\}:x_{\ell}\text{ has no removed edges}\}$. We define coupling $\pi$ as a probabilistic measure on $\mathbb{X}^{K+1}\times\mathbb{X}^{\prime}\times\mathbb{Y}^{2(K+1)}\times\mathbb{Y}^{\prime}$. As a shorthand we write $\vec{x}=(x_{0},...,x_{K})\in\mathbb{X}^{K+1}$, $\vec{y}^{(j)}=(y^{(j)}_{0},...,y^{(j)}_{K})\in\mathbb{Y}^{K+1}$ and $\vec{y}=(\vec{y}^{(0)},\vec{y}^{(1)})$. We require coupling $\pi(\vec{x},x^{\prime},\vec{y},y^{\prime})=\pi(\vec{x},x^{\prime})\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})$ to satisfy:
$\sum_{\vec{x}\backslash\{x_{i}\},x^{\prime}}\pi(\vec{x},x^{\prime})$ $=\omega(x_{i}|A_{i}),$ (8)
@@
Specifically, let us denote the size $|S(x_{i})|$ by $L$ and define $p_{n,m}\equiv 1/n(n-1)\cdots(n-m+1)$ be the inverse of falling factorial from $n$ to $n-m+1$, then we can explicitly write down the marginal probability:
@@
$=\sum_{i,j}\sum_{x_{i},y^{(j)}_{i}}\left(\sum_{\vec{x}\backslash\{x_{i}\},x^{\prime}}\pi(\vec{x},x^{\prime})\right)\left(\sum_{\vec{y}\backslash\{y^{(j)}_{i}\},y^{\prime}}\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})\right)\omega(A_{i})\omega(B_{j}|x_{i})\mu_{x_{i},y^{(j)}_{i}}$ (22)
@@
  • for each $i=1,2,...,K$, sample $x_{i}$ by randomly drawing a size-$i$ subset of the removed edges and adding the subset to $x_{0}$;
@@
  • for each $i=1,...,K$,
    • sample $y_{i}^{(0)}$ by randomly choosing $i$ nodes (not in $y^{(0)}_{0}$) to and adding them to $y^{(0)}_{0}$;
@@
Now let’s verify the coupling constraints Eq.(8) to (11). For Eq.(8), when $i=0$, we have $\sum_{\vec{x}\backslash\{x_{0}\},x^{\prime}}\pi(\vec{x},x^{\prime})=\pi(x_{0})=\omega(x_{0}|A_{0})$ as desired. When $i\geq 1$,
$\sum_{\vec{x}\backslash\{x_{i}\},x^{\prime}}\pi(\vec{x},x^{\prime})=\sum_{x_{0}}\pi(x_{0})\pi(x_{i}|x_{0})=\gamma^{L}(1-\gamma)^{m-K-L}\cdot\frac{1}{\binom{K}{i}}=\omega(x_{i}|A_{i}),$ (33)
@@
For Eqn.(10), when $i=j=0$, we have $\sum_{\vec{y}\backslash\{y_{0}^{(0)}\},y^{\prime}}\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})=\pi(y_{0}^{(0)})=\omega(y^{(0)}_{0}|B_{0},x_{0})$ by our construction. When $i\neq 0,j=0$, recall that $p_{n,m}\equiv 1/n(n-1)\cdots(n-m+1)$ is the inverse of falling factorial from $n$ to $n-m+1$, then
$\begin{split}\sum_{\vec{y}\backslash\{y_{i}^{(0)}\},y^{\prime}}\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})=\sum_{y_{0}^{(0)}}\pi(y_{0}^{(0)}|x_{0})\pi(y^{(0)}_{i}|y^{(0)}_{0},x_{i})&=p_{n-1,L}p_{n-L,i}=p_{n-1,L}p_{n-L,|x_{i}|-L}\\ &=\omega(y_{i}^{(0)}|B_{0},x_{i}).\end{split}$ (34)
@@
$\begin{split}\sum_{\vec{y}\backslash\{y_{i}^{(1)}\},y^{\prime}}\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})&=\sum_{y_{0}^{(0)}}\pi(y_{0}^{(0)}|x_{0})\pi(y_{i}^{(1)}|y_{0}^{(0)},x_{i})=(n-L)\cdot p_{n-1,L}\frac{1}{L}p_{n-L,i}\\ &=\frac{1}{L}p_{n-1,L+i-1}=\frac{1}{L}p_{n-1,|x_{i}|-1}=\omega(y_{i}^{(1)}|B_{1},x_{i}).\end{split}$ (35)
@@
[…]

### v3

@@ 4.2 Coupled Sampling and Its Privacy Amplification
[…]
Coupled sampling extends the existing privacy amplification framework by generalizing the sampling process to a multi-step, dependent procedure, where the flexible and potentially intricate dependencies between steps pose new analytical challenges. In the absence of such dependencies, that is, $\mathcal{B}^{(2)}\sim q(\cdot\mid D^{(2)})$, the contributions of each sampling step can be decomposed, allowing existing privacy amplification results to be adapted: in this decoupled case, the mini-batch $\mathcal{B}=\{(\mathcal{B}^{(1)}_{i},\mathcal{B}^{(2)}_{i})\}_{i=1,2,\ldots}$ can be viewed as a single-step sampling from the Cartesian product dataset $D^{(1)}\times D^{(2)}$. However, when dependencies are introduced, this separation no longer holds.
@@
  $w\leftarrow\text{a random end in }e_{i}^{+}$, and $e_{i,j}^{-}\leftarrow(w,v_{i,j})$ $\triangleright$ pair $w$ with the sampled node $v_{i,j}$
Return $\mathcal{B}=\{T_{1},T_{2},...,T_{b}\}=\{(e_{i}^{+},e_{i,1}^{-},...,e_{i,k_{\mathrm{neg}}}^{-})\}_{i=1,...,b}$
@@
Initialize Randomly drop edges from $\mathcal{E}$, so that maximal node degree is capped by $K$ and obtain $\bar{\mathcal{E}}$ (Daigavane et al., 2021, Algorithm 1). Let positive sampling rate $\gamma\leftarrow b/|\bar{\mathcal{E}}|$$\triangleright$ Node degree capping
@@
  positive edges $E^{+}$ $\leftarrow$ independently choosing each positive relation from $\bar{\mathcal{E}}$ with probability $\gamma$.
@@
  $\bar{\mathbf{g}}_{t}\leftarrow\text{FREQ-CLIP}(\mathcal{B}_{t},\{\mathbf{g}_{t}(T_{1}),\mathbf{g}_{t}(T_{2}),...\},C)$ $\triangleright$ Adaptive clipping Algorithm 1
@@
In the context of relational learning, cardinality-dependent sampling can be realized by choosing negative sampling methods that are independent of the contents of the sampled positive edges. For instance, we adopt Poisson subsampling as positive sampling, and adopt Algorithm 2 for negative sampling, which first randomly draws some nodes from the node set (independent of specific positive edges), and then constructs contrastive pairs by pairing one end of positive edges with the sampled nodes. Note that this pairing may occasionally produce negative edges that are actually true positives—a known issue in node pairing methods like in-batch sampling (You et al., 2020). However, this occurs with a negligible probability, especially given the sparsity of real-world graphs.
@@
where $\mathrm{Bin}(m,\gamma)$ stands for the Binomial distribution, effective sampling rate $\Gamma_{\ell}\coloneqq 1-(1-\gamma)^{K}(1-\frac{\ell\cdot k_{\mathrm{neg}}}{n})$ and $\Psi_{\alpha}(P\|Q)\coloneqq\mathbb{E}_{x\sim Q}(P(x)/Q(x))^{\alpha}$ for two distributions $P,Q$. The proof is deferred to Appendix B. $\Psi_{\alpha}$ here is a univariate integral and can be numerically calculated up to arbitrary precision (Mironov et al., 2019). Although Theorem 3 adopts $S^{(1)}$ as Poisson sampling and $S^{(2)}$ as sampling without replacement, the analysis and results can be extended to other combinations of sampling strategies. We leave these generalizations to the extended version of this work.
@@
Note that the bound Eq.(3) depends on the maximal node degree $K$. To mitigate the influence of high-degree nodes, in practice we cap each node’s degree at a desired constant $K$ by randomly sampling its neighbors, same as Algorithm 1 of (Daigavane et al., 2021). The effect of degree capping on utility is explored in Figure 3 in our later experiments.
@@ 5 Experimental Results
Per-iteration RDP Bounds. We evaluate the RDP bound $\varepsilon(\alpha)$ (Eq.(3)), shown in the left two subfigures of Figure 2, which characterizes the per-iteration privacy loss in DP-SGD. Our bound is compared against two baselines: (a) the naive Gaussian mechanism bound $\varepsilon_{\mathrm{Gaussian}}(\alpha)=\alpha/2\sigma^{2}$ (Mironov, 2017) without any amplification analyses; (b) the amplification bound with standard gradient clipping in place of the proposed adaptive clipping. The latter is also a coupled-sampling amplification bound we derived using an analysis similar to that of Theorem 3, and is provided in Appendix C.
DP Bounds over Composition. We also compare the accumulated privacy loss of applying multiple iterations of DP-SGD in terms of $(\varepsilon,\delta)$-DP, tracked by composition theorems of RDP (Mironov, 2017) and translating RDP to DP Balle et al. (2020). We choose privacy budget $\delta=1/m=0.2\times 10^{-6}$ and compare $\varepsilon_{\mathrm{DP}}(T)$ as a function of iteration $T$. The results of DP bound over composition are shown in the right two subfigures in Figure 2.
@@ 5.2 Private Relational Learning for Pretrained Text Encoders
Figure 3: Utility of different #\#negative sample per positive $k_{\text{neg}}$, degree capping $K$, and noise multiplier $\sigma$ for zero-shot relation prediction. The legend AMAZ-sports means training on AMAZ-cloth and testing on AMAZ-sports, and similar for the legend AMAZ-cloth. In the rightmost figure, numbers along each line indicate the corresponding privacy parameter $\varepsilon_{\mathrm{DP}}$ at different noise levels.
Experimental Settings. We adopt two pre-trained language models BERT (Devlin et al., 2019) and Llama2 (Touvron et al., 2023) as the text encoders, and four text-attributed graphs from two subdomain pairs: two citation networks (MAG-CHN, MAG-USA) (Sinha et al., 2015) and two co-purchase networks (AMZA-Sports, AMZA-Cloth) (McAuley et al., 2015). The models are first privately fine-tuned on one network (e.g., MAG-CHN), and then its utility is evaluated on another same-domain network (e.g., MAG-USA). We report the overall privacy loss in terms of $(\varepsilon,\delta)$-DP, where $\delta$ is set to $1/|\mathcal{E}_{\text{train}}|$, the size of the relation set used for training after degree capping. We use the ranking metrics of top@1 precision (PREC@1) and mean reciprocal rank (MRR) to evaluate the relation prediction performance. The maximal node degree is capped by $K=5$ according to Algorithm 3. See Appendix E for more implementation details.
@@ Appendix B Proof of Theorem 3
Notation. We use the removal/insertion notation of DP: we have two neighboring graphs $G=(\mathcal{V},\mathcal{E}),G^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ where $\mathcal{V}=\{1,2,...,n\}$, $\mathcal{V}^{\prime}=\{1,2,...,n-1\}$ and $\mathcal{E}=\mathcal{E}^{\prime}\cup\{(n,u_{i}):i=1,2,...,K\}$. Let $m=|\mathcal{E}|$. Let $\mathbb{X}=2^{\mathcal{E}},\mathbb{X}^{\prime}=2^{\mathcal{E}^{\prime}}$ be the all possible subset of edges that represents the space of positive samplings. Let $\mathbb{Y}=\{(y_{1},y_{2}...,y_{b\cdot k_{\mathrm{neg}}}):y_{i}\in V,b\in[m],y_{i}\neq y_{j}\}$ represent all possible ordered node set that represents the space of negative samplings. For simplicity, we will assume $k_{\mathrm{neg}}=1$, i.e., each positive sample is only paired with one negative sample. In the end we will show how to adapt to case $k_{\mathrm{neg}}>1$. An actual dataset we are computing on is $(x,y)$ for some $x\in\mathbb{X},y\in\mathbb{Y}$ or $(x^{\prime},y^{\prime})$ for some $x^{\prime}\in\mathbb{X}^{\prime},y\in\mathbb{Y}^{\prime}$ (again, $\mathbb{X}^{\prime},\mathbb{Y}^{\prime}$ are the sampling space of positive edges and ordered node set defined on neighboring graph $\mathcal{G}^{\prime}$). We denote $\mu_{x,y}$ as the output distribution of the mechanism acting on a fixed sampled dataset $(x,y)$, and denote $p,q$ as the mixture output distribution of the mechanism acting on a randomly sampled dataset $(x,y)$ and $(x^{\prime},y^{\prime})$ respectively. By definition, we have
@@
Jensen’s inequality by conditional coupling. The key idea is to apply Jensen’s inequality to $\Psi_{\alpha}(\cdot||\cdot)$ as $\Psi_{\alpha}$ is jointly convex [60]. To do so, we need to re-write $p,q$ in terms of balancing summation. we will generalize the conditional coupling technique [52] to our case. In the following we will show how we can achieve this in a high-level view. Consider the conditional distributions $\omega(x|A_{i})$, where $A_{i}$ represents the event “$x$ has $i$ removed edges (i.e., the edges involve the removed node nn)”, and conditional distributions $\omega(y|B_{j},x)$, $j=0,1$, where $B_{0}$($B_{1}$) represents the event $y]_{S(x)}$ has (no) removed node nn”, where $S(x)$ is a subset of indices defined by $S(x)\equiv\{\ell\in\{1,2,...,|x|\}:x_{\ell}\text{ has no removed edges}\}$. We define coupling $\pi$ as a probabilistic measure on $\mathbb{X}^{K+1}\times\mathbb{X}^{\prime}\times\mathbb{Y}^{2(K+1)}\times\mathbb{Y}^{\prime}$. As a shorthand we write $\vec{x}=(x_{0},...,x_{K})\in\mathbb{X}^{K+1}$, $\vec{y}^{(j)}=(y^{(j)}_{0},...,y^{(j)}_{K})\in\mathbb{Y}^{K+1}$ and $\vec{y}=(\vec{y}^{(0)},\vec{y}^{(1)})$. We require coupling $\pi(\vec{x},x^{\prime},\vec{y},y^{\prime})=\pi(\vec{x},x^{\prime})\pi(\vec{y},y^{\prime}|\vec{x},x^{\prime})$ to satisfy:
[…]
@@
Specifically, let us denote the size $|S(x_{i})|$ by $L$ and define $p_{n,m}\equiv 1/n(n-1)\cdots(n-m+1)$ be the inverse of falling factorial from $n$ to nm+1n-m+1, then we can explicitly write down the marginal probability:
@@
[…]
@@
  • for each i=1,2,,Ki=1,2,...,K, sample $x_{i}$ by randomly drawing a size-$i$ subset of the removed edges and adding the subset to $x_{0}$;
@@
  • for each i=1,,Ki=1,...,K,
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Derive explicit bound. Now we are going to derive bounds for Eqns.(28) and (29) by plugging $\pi$ into it. Note that coefficients $\{\omega(A_{i})\omega(B_{j}|x_{i})\}_{i,j}$ satisfy $\sum_{i,j}\omega(A_{i})\omega(B_{j}|x_{i})=1$ and thus $\sum_{i,j\neq 0}\omega(A_{i})\omega(B_{j}|x_{i})=1-\omega(A_{0})\omega(B_{0}|x_{0})$. Therefore, we can write